Please show me how to solve:
(100p-4000)/100p=.3

Found 2 solutions by Alan3354, scott8148:
Answer by Alan3354(69443)

Please show me how to solve:
(100p-4000)/100p=.3
------------
Multiply thru by 100p
(100p-4000) = 30p
70p = 4000
p = 400/7

Answer by scott8148(6628)

multiplying by 100p __ 100p-4000=30p __ adding 4000 __ 100p=30p+4000

subtracting 30p __ 70p=4000 __ dividing by 70 __ p=400/7